x-12=sq root (16x)
Answer by stanbon(48558)

x-12=sq root (16x)
Square both sides to get:
x^2-24x+144 = 16x
x^2-40x+144 = 0
(x-36)(x-4) = 0
x = 36 or x = 4
-----------------
Checking:
Does x=36 ?
36-12 = sqrt(16*36)
24 = 4*6
-------------
Does x = 4 ?
4-12 = sqrt(16*4)
-8 = 4*2
Wrong
----------------
Only solution: x= 36
========================
cheers,
stan H.